OpenStudy (anonymous):

find the polynomial function with roots 1, -2, and 5

OpenStudy (solomonzelman):

when you have a root of -2 it is written as (x+2) saying that when x=-2 then (x+2) is zero. So root of 1 would be written as (x+1) let say now that you have a function with roots of 3,4,-5 then you would write it as f(x)=(x-3)(x-4)(x+5) can you do your problem now, or still need more help ?

OpenStudy (anonymous):

So it would be f(x)=(x-1)(x+2)(x-5) ?

OpenStudy (solomonzelman):

yes, you are correct. I am not sure if you have to expand it or not, but I normally would.